# Clustering is back: Reaching state-of-the-art LiDAR instance segmentation without training

### v1

@@ Abstract
Panoptic segmentation of LiDAR point clouds is fundamental to outdoor scene understanding, with autonomous driving being a primary application. While state-of-the-art approaches typically rely on end-to-end deep learning architectures and extensive manual annotations of instances, the significant cost and time investment required for labeling large-scale point cloud datasets remains a major bottleneck in this field. In this work, we demonstrate that competitive panoptic segmentation can be achieved using only semantic labels, with instances predicted without any training or annotations. Our method achieves performance comparable to current state-of-the-art supervised methods on standard benchmarks including SemanticKITTI and nuScenes, and outperforms every publicly available method on SemanticKITTI as a drop-in instance head replacement, while running in real-time on a single-threaded CPU and requiring no instance labels. Our method is fully explainable, and requires no learning or parameter tuning. Code is available at https://github.com/valeoai/Alpine/
@@ 4 Experiments
LPSAD (in [51]) [39] 37.4 44.2 47.8 66.9 49.4
Panop.TrackNet [22] 40.0 - 48.3 73.0 53.8
PointGroup [23] 46.1 54.0 56.6 74.6 55.7
TORNADO-Net [17] 50.6 55.9 62.1 74.9 59.2
Panoster [16] 55.6 - 66.8 79.9 61.1
@@
Table 1: Panoptic segmentation results on the validation set of SemanticKITTI. ‘TTA’ and ‘Ens.’ stand for Test-Time Augmentation and Ensembling, respectively. (✓) denotes that test-time-augmentations and/or ensemble was used on the semantic segmentation head only. The main metric is the PQ.
@@
LPSAD (in [51]) [39] 50.4 57.7 62.4 79.4 62.5
Panop.TrackNet [22] 51.4 56.2 63.3 80.2 58.0
VIN [85] 51.7 57.4 61.8 82.6 73.7
TORNADO-Net [17] 54.0 59.8 65.4 80.9 68.0
MaskPLS [35] 57.7 60.2 66.0 71.8 62.5
GP-S3Net [51] 61.0 67.5 72.0 84.1 75.8
EfficientLPS [58] 62.0 65.6 73.9 83.4 65.6
@@
Table 2: Panoptic segmentation results on the validation set of nuScenes. ‘TTA’ and ‘Ens.’ stand for Test-Time Augmentation and Ensembling, respectively. (✓) denotes that test-time-augmentations and/or ensemble was used on the semantic segmentation head only. The main metric is the PQ.
@@ 4.4 Study of ALPINE’s components
Tab. 4 shows how all panoptic methods for which we found code and checkpoints on SemanticKITTI compare against ALPINE at equal semantic prediction: we use the semantic prediction of the original method but replace the instance prediction mechanism with ALPINE. We show that ALPINE outperforms every single method, while using no annotation and requiring no GPU. This further demonstrates that instance prediction does not need complex design, training or instance annotations to be competitive. An indicative “instance runtime” of all methods have been obtained by subtracting the segmentation forward pass to the total runtime, when this was possible. This approximate measurement is comparable (while not directly equivalent) to the ALPINE FPS reported in Tab. 5.
@@ 4.5 Evaluation of the clustering upper bound
Table 9: Comparison between statistics-based and web-based parameters settings. More run details can be found in Secs. 4 and 4.
@@ Appendix F Per-class results
For completeness, we present in Tabs. 11 and 12 the per-class results of our method on both nuScenes and SemanticKITTI.
Appendix G More visualisations
@@
Table 13: Panoptic segmentation results on the validation set of SemanticKITTI. ‘TTA’ and ‘Ens.’ stand for Test-Time Augmentation and Ensembling, respectively. (✓) denotes that test-time-augmentations and/or ensemble was used on the semantic segmentation head only.
@@
Table 14: Panoptic segmentation results on the validation set of nuScenes. ‘TTA’ and ‘Ens.’ stand for Test-Time Augmentation and Ensembling, respectively. (✓) denotes that test-time-augmentations and/or ensemble was used on the semantic segmentation head only.

### v2

@@ Abstract
Panoptic segmentation of LiDAR point clouds is fundamental to outdoor scene understanding, with autonomous driving being a primary application. While state-of-the-art approaches typically rely on end-to-end deep learning architectures and extensive manual annotations of instances, the significant cost and time investment required for labeling large-scale point cloud datasets remains a major bottleneck in this field. In this work, we demonstrate that competitive panoptic segmentation can be achieved using only semantic labels, with instances predicted without any training or annotations. Our method outperforms state-of-the-art supervised methods on standard benchmarks including SemanticKITTI and nuScenes, and outperforms every publicly available method on SemanticKITTI as a drop-in instance head replacement, while running in real-time on a single-threaded CPU and requiring no instance labels. It is fully explainable, and requires no learning or parameter tuning. ALPINE combined with state-of-the-art semantic segmentation ranks first on the official panoptic segmentation leaderboard of SemanticKITTI. Code is available at https://github.com/valeoai/Alpine/
@@ 4 Experiments
Table 1: Panoptic segmentation results on the validation set of SemanticKITTI. ‘TTA’ and ‘Ens.’ stand for Test-Time Augmentation and Ensembling. (✓) denotes that TTA/ensemble was used on the semantic head only. The main metric is the PQ. A more detailed version of this table is available in the supplementary material.
@@
Table 2: Panoptic segmentation results on the validation set of nuScenes. ‘TTA’ and ‘Ens.’ stand for Test-Time Augmentation and Ensembling. (✓) denotes that TTA/ensemble was used on the semantic head only. The main metric is the PQ. A more detailed version of this table is available in the supplementary material.
@@ 4.3 Comparison to the state-of-the-art
[…]
5/ P3Former 65.3 67.8 74.9 86.6 67.4 66.1
4/ PUPS 65.7 70.3 75.8 85.7 68.1 68.5
3/ P3Former & ALPINE 66.1 68.7 75.9 86.7 69.4 66.1
2/ UniSeg 67.2 72.1 78.1 85.5 67.5 73.8
1/ UniSeg & ALPINE 70.2 75.1 80.5 86.7 74.3 75.1
Table 4: SemanticKITTI test set results (official leaderboad on June 12, 2025). We used P3Former’s and UniSeg’s semantic predictions and used ALPINE for instance segmentation.
@@ 4.4 Study of ALPINE’s components
We present results obtained with ALPINE on the test set of SemanticKITTI. First, we use the public model of P3Former for semantic prediction and extract object instances using ALPINE. The PQ which is at 65.3 with P3Former improves to 66.1 thanks to ALPINE, the third best entry on the leaderboard. Second, we repeat the same procedure with the semantic labels of UniSeg [35] (labels kindly provided by the authors). UniSeg with ALPINE’s instance labels ranks first on the leaderboard of SemanticKITTI. These results highlight the high quality of the instances extracted by our method. We note nevertheless that these results are not strictly obtained without instance labels as both UniSeg and P3Former did use them during training. We also remark that while the semantic mIoU of P3Former and P3Former&ALPINE are identical, UniSeg&ALPINE’s semantic mIoU is above UniSeg’s: 75.1 vs 73.8. We did not modify UniSeg’s semantic labels (75.1 is in line with the score of UniSeg on the semantic segmentation leaderboard). We suspect a misformatting in UniSeg’s panoptic labels but which we cannot correct ourselves.
We conduct a more complete study on the validation set of SemanticKITTI. Tab. 5 shows how all panoptic methods for which we found code and checkpoints on SemanticKITTI compare against ALPINE at equal semantic prediction: we use the semantic prediction of the original method but replace the instance prediction mechanism with ALPINE. We show that ALPINE outperforms every single method, while using no annotation and requiring no GPU. This further demonstrates that instance prediction does not need complex design, training or instance annotations to be competitive. An indicative “instance runtime” of all methods have been obtained by subtracting the segmentation forward pass to the total runtime, when this was possible. This approximate measurement is comparable (while not directly equivalent) to the ALPINE FPS reported in Tab. 6.
@@ 4.5 Evaluation of the clustering upper bound
Table 10: Comparison between statistics-based and web-based parameters settings. Run details can be found in Secs. 4 and 4.
@@ 6 Conclusion
Acknowledgements.
We wish to thank the authors of P3Former [71] for making their code and model used on the test set of SemanticKITTI available. We also thank the authors of UniSeg [35], particularly Lingdong Kong, for providing us with their test set semantic predictions. This work was supported in part by the French Agence Nationale de la Recherche (ANR) grant MultiTrans (ANR21-CE23-0032).
@@ Appendix F Per-class results
For completeness, we present in Tabs. 12 and 13 the per-class results of our method on both nuScenes and SemanticKITTI.
Appendix G Performance of ALPINE and distance to the sensor
0-15m 15-30m 30m+
[…]
ALPINE w/ Ens. 68.5 71.0 73.0 68.6 70.7 68.4 65.8 65.9 54.4
P3Former 63.6 65.8 67.8 64.8 66.9 63.4 59.6 60.1 47.2
P3Former & ALPINE 64.0 66.2 67.8 64.8 66.9 63.4 59.4 59.8 47.2
Table 14: Performance w.r.t sensor distance on SemanticKITTI’s validation set.
We report in Tab. 14 a breakdown of ALPINE’s performance with respect to the distance to the sensor. ALPINE improves over P3Former at short range, but slightly deteriorates at long range. This suggests that ALPINE is particularly efficient on dense point clouds (i.e. close to the sensor) but gets lesser performance on sparser point clouds.
Appendix H More visualisations
@@
Table 15: Panoptic segmentation results on the validation set of SemanticKITTI. ‘TTA’ and ‘Ens.’ stand for Test-Time Augmentation and Ensembling. (✓) denotes that TTA/ensemble was used on the semantic head only. The main metric is the PQ.
@@
Table 16: Panoptic segmentation results on the validation set of nuScenes. ‘TTA’ and ‘Ens.’ stand for Test-Time Augmentation and Ensembling. (✓) denotes that TTA/ensemble was used on the semantic head only. The main metric is the PQ.